# Channel Estimation for RIS-Assisted mmWave Systems via Diffusion Models

### v2

@@ Abstract
Index Terms:
Diffusion models, channel estimation, reconfigurable intelligent surface.
@@ I Introduction
Millimeter-wave (mmWave) communication is a key technology for the sixth generation (6G) communication systems to fully exploit the available spectrum [1]. Nevertheless, the coverage range of mmWave signals is limited due to their rapid attenuation, which makes them susceptible to blockage. To overcome this limitation, reconfigurable intelligent surfaces (RISs) [2] have been introduced to extend communication coverage and improve transmission rates. However, the effectiveness of the RIS depends on accurate channel state information (CSI). In general, a RIS consists of passive elements without signal processing capabilities, which poses a significant challenge for channel estimation in RIS-assisted systems.
@@
Conventional methods (e.g., linear minimum mean square error, LMMSE) and compressed sensing (CS)-based methods (e.g., the double-structured orthogonal matching pursuit, DS-OMP [3]) struggle to achieve accurate channel estimation under low signal-to-noise ratio (SNR) conditions. With the rapid development of artificial intelligence (AI), deep learning (DL)-based methods have been increasingly applied in RIS-assisted systems. For instance, a convolutional neural network (CNN)-based method is proposed in [4], where a deep residual network (DRN) is used to denoise the results obtained from the least squares (LS) estimator. Furthermore, a global attention residual network (GARN) is introduced to fuse multi-channel information and improve the accuracy of channel estimation [5]. However, these direct-mapping DL methods require network retraining for different SNRs, limiting their generalizability and practicality.
Recently, diffusion models (DMs) [6] have emerged as powerful generative AI techniques capable of modeling complex data distributions. To leverage their strengths, various studies have explored the application of DMs in wireless communications. Specifically, channel denoising diffusion models (CDDMs) are proposed in [7] as a separate module following the channel equalizer. In [8], DMs are integrated with low-density parity-check (LDPC) decoding to enhance error correction. Moreover, DMs are applied to high-dimensional channel estimation in [9]. Among these applications, channel estimation has garnered particular attention due to its critical role in communication systems. In [10], guided DMs are employed for channel estimation, yet the U-Net architecture introduces significant memory and computational overhead. To address this issue, a CNN-based network that achieves lower overhead is proposed in [11]. Nevertheless, this simplified architecture suffers from performance degradation in more complex propagation environments. Therefore, achieving accurate and efficient channel estimation in RIS-assisted systems remains a challenging problem that requires further investigation.
@@
In this letter, we propose a novel channel estimation method for RIS-assisted mmWave systems based on DMs. The key contributions are summarized as follows
@@
  • We formulate the channel estimation problem in RIS-assisted systems as the reverse denoising process of denoising diffusion implicit models (DDIMs). This approach inherently adapts to dynamic noise conditions through a step-matching mechanism.
@@
  • We first propose a U-Net-based channel estimation scheme and subsequently develop a parameter-efficient network architecture, referred to as BRCNet, to reduce model complexity and computational overhead.
@@
  • Extensive simulations demonstrate that the proposed method outperforms the baselines across various scenarios, while the BRCNet achieves performance comparable to U-Net using only 9.34% of its parameters.
@@ II-A Channel Model
As illustrated in Fig. 1, we consider the uplink channel estimation for RIS-assisted mmWave systems. Both the base station (BS) and RIS are equipped with uniform planar arrays (UPAs) with half-wavelength element spacing, while the user equipment (UE) has a single antenna. The BS has $M=M_{h}\times M_{v}$ antennas, and the RIS consists of $N=N_{h}\times N_{v}$ passive elements. The UE-BS, UE-RIS, and RIS-BS channels are denoted as $\mathbf{d}\in\mathbb{C}^{M\times 1}$, $\mathbf{h}\in\mathbb{C}^{N\times 1}$, and $\mathbf{G}\in\mathbb{C}^{M\times N}$, respectively. Following [12] and [13], a clustered statistical channel model is adopted. The RIS-BS channel consists of line-of-sight (LOS) and non-line-of-sight (NLOS) components, i.e., $\mathbf{G}=\mathbf{G}_{\text{LOS}}+\mathbf{G}_{\text{NLOS}}$. These two components are represented as
@@
×$\times\mathbf{a}^{T}_{\text{RIS}}\left(\phi^{G_{t}}_{\text{LOS}},\varphi^{G_{t}}_{\text{LOS}}\right),$
@@
×$\times\mathbf{a}_{\text{RIS}}^{T}\left(\phi^{G_{t}}_{c,s},\varphi^{G_{t}}_{c,s}\right),$
where $C^{G}$ and $S^{G}_{c}$ denote the number of clusters and the number of scatterers in the $c$-th cluster, respectively. The normalization factor is defined as $\gamma^{G}=\sqrt{\frac{1}{\sum_{c=1}^{C}S_{c}}}$. $\beta^{G}_{c,s}\sim\mathcal{CN}(0,1)$ denotes the path gain of the $(c,s)$-th scatterer, which follows a circularly symmetric complex Gaussian (CSCG) distribution. The element radiation pattern of the RIS corresponding to the $(c,s)$-th scatterer is defined as $G_{e}\left(\varphi^{G_{t}}_{c,s}\right)=2(2\zeta+1)\cos^{2\zeta}(\varphi^{G_{t}}_{c,s})$ [14]. The path loss $L_{c,s}^{G_{t}}$ is defined as
$L_{c,s}^{G_{t}}=$ $-20\log_{10}\left(\frac{4\pi}{\lambda}\right)$ (3)
$-10n\left(1+b\left(\frac{f_{c}-f_{0}}{f_{0}}\right)\right)\log_{10}(d_{c,s})-X_{\sigma_{x}},$
where $\lambda$ is the carrier wavelength, $n$ is the path loss exponent, $b$ is a system parameter, $f_{0}$ is the reference frequency, $f_{c}$ is the carrier frequency, $d_{c,s}$ is the propagation distance of the $(c,s)$-th scatterer, and $X_{\sigma_{x}}\sim\mathcal{CN}(0,\sigma_{x}^{2})$ is the shadow factor.
@@
$\mathbf{a}(\phi,\varphi)=[$ $1,\ldots,e^{j2\pi d(x\sin\varphi+y\sin\phi\cos\varphi)/\lambda},$ (4)
$\ldots,e^{j2\pi d\left((N_{h}-1)\sin\varphi+(N_{v}-1)\sin\phi\cos\varphi\right)/\lambda}]^{T},$
where $10\leq x\leq N_{h}-$ 1 and $10\leq y\leq N_{v}-$ 1, and $d$ denotes the element spacing. Similarly, the UE-RIS channel $\mathbf{h}$ and the UE-BS channel $\mathbf{d}$ share the same structural form as $\mathbf{G}$.
@@ II-B Signal Model
where $\mathbf{H}=\left[\mathbf{d},\ \mathbf{H}_{c}\right]\in\mathbb{C}^{M\times(N+1)}$. Without loss of generality, the pilot symbol is set to $x_{k}=1$ [5, 10]. After $K$ time slots, the received signal $\mathbf{Y}\in\mathbb{C}^{M\times K}$ is given by
@@
The RIS reflecting matrix is set as the columns of the discrete Fourier transform (DFT) matrix [15]. The objective of channel estimation is to recover the unknown channel $\mathbf{H}$ from the received signal $\mathbf{Y}$. By vectorizing (7), we obtain
@@
where $\text{vec}(\cdot)$ denotes vectorization and $\otimes$ denotes the Kronecker product. Given the noise power $\sigma^{2}$, (8) is normalized and converted into an equivalent real-valued form as
@@ III Proposed Method
This section details the proposed method. First, a forward diffusion process for the received signal is formulated to define the training algorithm. Then, a sampling algorithm based on the DDIM is developed. Furthermore, a lightweight neural network architecture, termed BRCNet, is designed to reduce the model complexity and computational overhead.
@@ III-A Training Algorithm of the Proposed Method
The original signal $\mathbf{x}_{0}$ is defined as $\mathbf{x}_{0}=\mathbf{A}\mathbf{p}$. By progressively adding noise to $\mathbf{x}_{0}$, the signal at step $t\in\{1,2,...,T\}$ is obtained as follows
$\mathbf{x}_{t}=\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}+\beta_{t}\bm{\epsilon}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},$ (10)
@@
The proposed method aims to learn the distribution of $\mathbf{x}_{0}$ for effective denoising. This can be achieved by optimizing the variational bound on negative log-likelihood, given by [6]
@@
where $D_{\text{KL}}(\cdot)$ denotes the Kullback-Leibler (KL) divergence. The third term measures the mismatch between the true distribution and the parameterized distribution and is optimized to ensure that the parameterized distribution closely approximates the true one. $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ can be derived via Bayes’ rule as
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\frac{q(\mathbf{x}_{t}|\mathbf{x}_{t-1},\mathbf{x}_{0})q(\mathbf{x}_{t-1}|\mathbf{x}_{0})}{q(\mathbf{x}_{t}|\mathbf{x}_{0})}.$ (13)
Combining (11) and (13), we derive
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ (14)
$\sim\mathcal{N}\left(\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\mathbf{x}_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\mathbf{x}_{t},\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}\mathbf{I}\right).$
With a specific parameterization of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, defined as
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ (15)
$\sim\mathcal{N}\left(\mathbf{x}_{t-1};\ \frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right),\ \sigma_{t}^{2}\mathbf{I}\right),$
$L_{t-1}$ can be calculated in a Rao-Blackwellized fashion with closed-form expressions [6]. To summarize, the neural network’s loss function can be formulated as
@@
Through the training process, the proposed method is capable of estimating the noise in $\mathbf{x}_{t}$, as $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is optimized to approximate $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$. To facilitate training, we adopt the angular domain representation as in [11]. Accordingly, the training dataset is constructed as $\mathcal{X}=\left\{\mathbf{X}_{n}=\operatorname{fft}(\mathbf{H}_{n}\mathbf{P})\right\}_{n=1}^{N_{\mathrm{train}}}$, where $\operatorname{fft}(\cdot)$ denotes the two-dimensional (2D) fast Fourier transform (FFT). In summary, the training algorithm is outlined in Algorithm $1$Consistent with [6] and [11], the vectorized form is employed for theoretical analysis, whereas the matrix form is used in practical implementations..
@@
0:  $\hat{\mathbf{H}}$
1:  $\mathbf{X}_{t_{s}}=\operatorname{fft}(\frac{1}{\sqrt{1+\sigma^{2}}}\mathbf{Y})$;
@@
6:  $\mathbf{X}_{0}=\frac{1}{\sqrt{\alpha_{1}}}\left(\mathbf{X}_{1}-\sqrt{1-\bar{\alpha}_{1}}\bm{\epsilon}_{\theta}(\mathbf{X}_{1},1)\right)$;
7:  $\hat{\mathbf{H}}=\operatorname{ifft}(\mathbf{X}_{0})\mathbf{P}^{\dagger}$;
@@ III-B Sampling Algorithm of the Proposed Method
Based on the trained network $\bm{\epsilon}_{\theta}$, an estimate of $\mathbf{x}_{0}$ can be obtained by replacing $\bm{\epsilon}$ with $\bm{\epsilon}_{\theta}$ in (10), as follows
$\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right).$ (17)
According to [16], $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ can be estimated from $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{x}}_{0})$. Therefore, the sampling process is given by
$\mathbf{x}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\underbrace{\left(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)}_{\text{estimate }\mathbf{x}_{0}}$ (18)
$+\underbrace{\sqrt{1-\bar{\alpha}_{t-1}}\,\epsilon_{\theta}(\mathbf{x}_{t},t)}_{\text{direction pointing to }\mathbf{x}_{t}}.$
Given (9) and (10), when $\sqrt{\bar{\alpha}_{t_{s}}}=\frac{1}{\sqrt{1+\sigma^{2}}}$, the KL divergence is $D_{\text{KL}}(q(\mathbf{x}_{t_{s}}|\mathbf{x}_{0})\parallel p(\mathbf{y}|\mathbf{x}_{0}))=0$. This indicates that the forward process yields signals that share the same distribution as the received signals. Here, $\mathbf{x}_{t_{s}}$ denotes the signal at step $t_{s}$ in the forward process. Therefore, during the sampling process, we can start from an intermediate step $t_{s}$ rather than from step $T$ as done in [6], which significantly improves the efficiency. The index $t_{s}$ is computed as
[…]
which is referred to as the step-matching mechanism. The starting step depends on the noise level, with larger indices indicating stronger noise contamination and thus requiring more denoising steps. By incorporating (19), the proposed approach can adapt to varying noise conditions without retraining.

### v3

@@ I Introduction
Millimeter-wave (mmWave) communication is a key technology for the sixth generation (6G) communication systems to fully exploit the available spectrum [Wang2023]. Nevertheless, the coverage range of mmWave signals is limited due to their rapid attenuation, which makes them susceptible to blockage. To overcome this limitation, reconfigurable intelligent surfaces (RISs) [Wu2019] have been introduced to extend communication coverage and improve transmission rates. However, the effectiveness of the RIS depends on the acquisition of accurate channel state information (CSI). In general, a RIS consists of passive elements without signal processing capabilities, which poses a significant challenge for channel estimation in RIS-assisted systems.
@@
Conventional methods (e.g., linear minimum mean square error, LMMSE) and compressed sensing (CS)-based methods (e.g., the double-structured orthogonal matching pursuit, DS-OMP [Wei2021]) struggle to achieve accurate channel estimation under low signal-to-noise ratio (SNR) conditions. With the rapid development of deep learning (DL), DL-based methods have been increasingly applied in RIS-assisted systems. For instance, a deep residual network (DRN) is used to denoise the results obtained from the least squares (LS) estimator in [Liu2021]. Furthermore, a global attention residual network (GARN) is introduced to fuse multi-channel information and improve the accuracy of channel estimation [Feng2023]. However, these direct-mapping DL methods require network retraining for different SNRs, limiting their generalizability and practicality.
Recently, diffusion models (DMs) [Ho2020] have garnered significant attention in wireless communications due to their strong capability to model complex data distributions [Wu2024, Guan2024, Zhou2025, Fesl2024]. Among their various applications, channel estimation has received particular attention due to its critical role, especially in RIS-assisted scenarios where cascaded and direct links lead to more complex distributions and greater challenges. In [Tong2024], the authors employ received signals to guide the sampling of DMs for channel estimation. Another approach has been proposed in [Zhang2025], which infers the full channel from the partial channel based on conditional DMs. Moreover, in [Liu2022], the diffusion process is conducted in the spectral domain to improve stability. However, in these works, the sampling process typically starts from the maximum step $T$ without adapting to different signal-to-noise ratio (SNR) conditions, which limits their robustness under varying channel environments. Furthermore, the network architectures adopted in previous works fail to strike a balance among performance, parameter size, and computational complexity. Therefore, achieving accurate and efficient channel estimation in RIS-assisted systems remains a challenging problem that requires further exploration.
@@
In this letter, we propose a novel and practical channel estimation method for RIS-assisted mmWave systems based on DMs. The key contributions can be summarized as follows.
@@
  • We formulate both direct and cascaded channel estimation in RIS-assisted scenarios within the framework of denoising diffusion implicit models (DDIMs), and introduce a step-matching mechanism that allows the estimation process to adapt to different SNR conditions.
@@
  • We initially adopt a U-Net-based channel estimation scheme as a baseline, and subsequently propose a parameter-efficient architecture, termed BRCNet, achieving an improved trade-off among estimation performance, model complexity, and computational overhead.
@@
  • Extensive simulations demonstrate that the proposed method outperforms the baselines under various scenarios, while the proposed BRCNet achieves performance comparable to U-Net using only 9.34% of its parameters.
@@ II-A Channel Model
As illustrated in Fig. 1, we consider the uplink channel estimation for RIS-assisted mmWave systems. Both the base station (BS) and RIS are equipped with uniform planar arrays (UPAs) with half-wavelength element spacing, while the user equipment (UE) has a single antenna. The BS has $M=M_{h}\times M_{v}$ antennas, and the RIS consists of $N=N_{h}\times N_{v}$ passive elements. The UE-BS, UE-RIS, and RIS-BS channels are denoted as $\mathbf{d}\in\mathbb{C}^{M\times 1}$, $\mathbf{h}\in\mathbb{C}^{N\times 1}$, and $\mathbf{G}\in\mathbb{C}^{M\times N}$, respectively. Following [3GPP2018] and [Basar2021], a clustered statistical channel model is adopted. The RIS-BS channel consists of line-of-sight (LOS) and non-line-of-sight (NLOS) components, i.e., $\mathbf{G}=\mathbf{G}_{\text{LOS}}+\mathbf{G}_{\text{NLOS}}$. These two components are represented as
@@
$\times\mathbf{a}^{T}_{\text{RIS}}\left(\phi^{G_{t}}_{\text{LOS}},\varphi^{G_{t}}_{\text{LOS}}\right),$
@@
$\times\mathbf{a}_{\text{RIS}}^{T}\left(\phi^{G_{t}}_{c,s},\varphi^{G_{t}}_{c,s}\right),$
where $C^{G}$ and $S^{G}_{c}$ denote the number of clusters and the number of scatterers in the $c$-th cluster, respectively. The normalization factor is defined as $\gamma^{G}=\sqrt{\frac{1}{\sum_{c=1}^{C^{G}}S_{c}^{G}}}$. $\beta^{G}_{c,s}\sim\mathcal{CN}(0,1)$ denotes the path gain of the $(c,s)$-th scatterer, which follows a circularly symmetric complex Gaussian (CSCG) distribution. The element radiation pattern of the RIS corresponding to the $(c,s)$-th scatterer is defined as $G_{e}\left(\varphi^{G_{t}}_{c,s}\right)=2(2\zeta+1)\cos^{2\zeta}(\varphi^{G_{t}}_{c,s})$. The path loss $L_{c,s}^{G_{t}}$ is defined as
[…]
where $\lambda$ is the carrier wavelength, $n$ is the path loss exponent, $b$ is a system parameter, $f_{0}$ is the reference frequency, $f_{c}$ is the carrier frequency, $d_{c,s}$ is the propagation distance of the $(c,s)$-th scatterer, and $X_{\sigma_{x}}\sim\mathcal{N}(0,\sigma_{x}^{2})$ is the shadow factor.
@@
$\mathbf{a}(\phi,\varphi)=[$ $1,\ldots,e^{j\frac{2\pi d}{\lambda}(x\sin\varphi+y\sin\phi\cos\varphi)},$ (4)
$\ldots,e^{j\frac{2\pi d}{\lambda}\left((N_{h}-1)\sin\varphi+(N_{v}-1)\sin\phi\cos\varphi\right)}]^{T},$
where $0\leq x\leq N_{h}-1$ and $0\leq y\leq N_{v}-1$, and $d$ denotes the element spacing. Similarly, the UE-RIS channel $\mathbf{h}$ and the UE-BS channel $\mathbf{d}$ share the same structural form as $\mathbf{G}$.
@@ II-B Signal Model
where $\mathbf{H}=\left[\mathbf{d},\ \mathbf{H}_{c}\right]\in\mathbb{C}^{M\times(N+1)}$. Without loss of generality, the pilot symbol is set to $x_{k}=1$ [Feng2023, Tong2024]. After $K$ time slots, the received signal $\mathbf{Y}\in\mathbb{C}^{M\times K}$ is given by
@@
The RIS reflecting matrix is set as the columns of the discrete Fourier transform (DFT) matrix [Zheng2020]. The objective of channel estimation is to recover the unknown channel $\mathbf{H}$ from the received signal $\mathbf{Y}$. By vectorizing (7), we obtain
@@
where $\operatorname{vec}(\cdot)$ denotes vectorization and $\otimes$ denotes the Kronecker product. Given the noise power $\sigma^{2}$, (8) is normalized and converted into an equivalent real-valued form as
@@ III Proposed Method
This section details the proposed method. First, a forward diffusion process for the received signal is formulated to define the training procedure. Then, a sampling algorithm based on DDIMs is developed. Furthermore, a lightweight neural network architecture, termed BRCNet, is designed to reduce model complexity and computational overhead.
@@ III-A Training Algorithm of the Proposed Method
For clarity, the original signal in the diffusion process is defined as $\mathbf{x}_{0}=\mathbf{A}\mathbf{p}$. By progressively adding noise to $\mathbf{x}_{0}$, the forward diffusion process simulates the effect of varying noise levels on the pilot signal after RIS-assisted transmission. Consequently, the signal at step $t\in\{1,2,...,T\}$ is given by
$\mathbf{x}_{t}=\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}+\sqrt{\beta_{t}}\bm{\epsilon}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},$ (10)
@@
The objective of the diffusion model is to learn the distribution of $\mathbf{x}_{0}$, enabling accurate reconstruction of clean pilot responses from noisy observations. Following the DMs framework [Ho2020], minimizing the variational bound leads to a noise-prediction objective, which is given by
@@
where $D_{\text{KL}}(\cdot)$ denotes the Kullback-Leibler (KL) divergence. The third term captures the gap between the true backward transition and its parameterized approximation. By aligning the parameterized transition with the true conditional distribution of the noisy signal, the model can achieve more accurate denoising and signal reconstruction in the RIS-assisted channel scenario. $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ can be derived via Bayes’ rule as in [Ho2020]. With a specific parameterization of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, $L_{t-1}$ can be calculated in a Rao-Blackwellized fashion with closed-form expressions [Ho2020]. To summarize, the neural network’s loss function can be formulated as
@@
By optimizing (13), the network learns to estimate the noise component in $\mathbf{x}_{t}$, enabling progressive reconstruction of the clean signal during sampling. The inherent angular sparsity of RIS-assisted mmWave channels aligns well with the progressive denoising nature of diffusion models, facilitating more effective learning of the channel structure. Therefore, we adopt the angular-domain representation as in [Fesl2024], and construct the training dataset as $\mathcal{X}=\left\{\mathbf{X}_{n}=\operatorname{fft}(\mathbf{H}_{n}\mathbf{P})\right\}_{n=1}^{N_{\mathrm{train}}}$, where $\operatorname{fft}(\cdot)$ denotes the two-dimensional fast Fourier transform (FFT). In summary, the training algorithm is outlined in Algorithm 1111Consistent with [Ho2020] and [Fesl2024], the vectorized form is employed for theoretical analysis, whereas the matrix form is used in practical implementations..
@@
[…]
@@
[…]
@@ III-B Sampling Algorithm of the Proposed Method
During the sampling process, the proposed method progressively denoises the signal through the sequential transitions $\mathbf{x}_{t}\rightarrow\mathbf{x}_{t-1}\rightarrow\cdots\rightarrow\mathbf{x}_{0}$. This stepwise refinement restores the dominant angular components of the RIS-assisted channel from noisy observations, leading to more accurate channel estimation. According to [Song2021], $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ is approximated by $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{x}}_{0})$, yielding the following sampling process
$\hat{\mathbf{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\underbrace{\left(\frac{\hat{\mathbf{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}(\hat{\mathbf{x}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)}_{\text{estimate of }\mathbf{x}_{0}}$ (14)
$+\underbrace{\sqrt{1-\bar{\alpha}_{t-1}}\,\epsilon_{\theta}(\hat{\mathbf{x}}_{t},t)}_{\text{direction pointing to }\mathbf{x}_{t}}.$
Given (9) and (10), when $\sqrt{\bar{\alpha}_{t_{s}}}=\frac{1}{\sqrt{1+\sigma^{2}}}$, the KL divergence satisfies $D_{\mathrm{KL}}(q(\mathbf{x}_{t_{s}}|\mathbf{x}_{0})\parallel p(\mathbf{y}|\mathbf{x}_{0}))=0$. This indicates that the forward process yields signals that follow the same distribution as the received signals. Here, $\mathbf{x}_{t_{s}}$ denotes the signal at step $t_{s}$ in the forward process. Therefore, during the sampling process, we can start from an intermediate step $t_{s}$ rather than from step $T$ as done in [Ho2020], which significantly improves the efficiency. The index $t_{s}$ is computed as
$t_{s}=\underset{t}{\arg\min}\left|2\sigma^{2}-\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}\right|,$ (15)
which is referred to as the step-matching mechanism [Wu2024]. The starting step depends on the noise level, with larger indices indicating stronger noise contamination and thus requiring more denoising steps. By incorporating (15), the proposed approach can adapt to varying noise conditions without retraining.